The rate of diffusion of 2 gases A and B are in the ration 16:3 If the ration of their masses present in the misture is 2:3 Then
 
a,. Then ratio o their molar masses is 16:1
b.. The ratio of thier molar masses is 1:4 
c the ratio of their moles present inside the container is 1:24
d the ratio of their moles present inside the container is 8:3

Grade:11

1 Answers

Vikas TU
7 years ago
We know the rate of diffusion formulae:
Therefore,
R1/R2 => 16/3 = (2x/M1)A/M1root(M1T)     *  M2root(M2T)/3x
we get M2/M1 = 64^(1/3) = 4/1
and hence,
M1/M2 = 1:4 is the answer.
where 2x/M1 = moles for A gas 
and
3x/M2 = moles for B gas respectively.
And we have placed them for pressure as from ideal gas eqn. Pressure is proprotional to the moles of the gases keeping the temperature constant.
Hence
option b.